4. (6 pts) Besides the Fed, the federal government also wants to counteract the effects of the virus outbreak using fiscal policy.

a. What are the two main policy steps the government can take?

b. Suppose the marginal propensity to consume is 0.75. How much will total consumption change if the government increases its spending by $1 billion?

c. How much will total consumption change if the government decreases taxes by $1 billion instead?

d. Why does cutting taxes have a different overall effect?

e. BONUS: Do you know what step the government actually considering in order to aid people and the economy? How would you classify this policy? (Hint: besides the two policies described above, there is also a third option we did not talk about in detail)
Answer #1

Fiscal policy tries to affect the economy by change in government spending or tax rate.

a) Government can either raise the government spending for treating the patients suffering from virus, raising awareness, providing food in case of lock down.

Or

Government can also reduce the tax rate such that people will get a hike in their disposable income and induced to spend more money in the economy.

b) MPC = 0.75

If government spending raises by $1 billion.

Multiplier = 1 / (1 - MPC)

Multiplier = 1 / (1 - 0.75) = 4

It means that if government raise expenditure by $1 billion, there would be a hike of $4 billion in GDP.

c) Tax Multiplier = MPC / (1 - MPC)

Tax Multiplier = 0.75 / (1 - 0.75) = 3

We can say that by reduction in taxes by $1 billion, there would be rsie in consumption to by $3 billion as people have high disposable income.

d) Cutting taxes would raise the disposable income of consumers which will raise the aggregate demand as demand and GDP have a positive relationship with each other.

Cutting taxes will raise incentive for producers to work more which will raise the overall productivity level.

e) There can be various reasons but one of them could be that government can lower the rate of interest such that people do not save their money and spend that money such that there is less reduction in circulation of money which becomes the main reason of lower economic growth.